Operations Report – February 20th

Crew 205 Operations Report 20 – FEB – 2019

SOL: 10

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1400 (20thFeb), turned off at 0800 (20thFeb)

Solar— SOC % – Turned off Gen (20th Feb) 57%, Turned on Gen (20th Feb) 90%

Diesel Reading – 75%

Propane Reading – ~ 34%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 39%, ~ 214 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 189 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01414004 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 49%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.3 hours

Beginning charge: 100%

Ending charge: 71%

Currently charging: Yes

Spirit rover used: Used

Hours: 76.1

Beginning charge: 100%

Ending charge: 53%

Currently charging: Yes

Notes on rovers: Spirit and Opportunity were not at 100% charge at the beginning of the first attempt at the EVA due to problems with water and the external breaker of RAM. Breaker was reset and is now functional. They were left to charge for an hour in order to reach 100% charge and were then fit to be used on EVA.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Journalist Report – February 20th

Crew 205 Crew Journalist Report 20-FEB-2019 Sol 10

Author: Maria Grulich

Title: “Moonshine and Puzzle Fever”

Crew 205 decided to start a new project: The Mars Puzzle! It seems to be tradition that crews finish this puzzle while being at MDRS and we are planning not to be an exception!

Since yesterday the fever has spread among the crew and every free minute not spent working is used to finish the 1000 piece puzzle. We will not stop until we have finished it and we will keep you posted!

The crew journalist could not resist grabbing her camera last night and shot pictures from the tunnels around the Hab with the Moon lighting up the sky almost like daytime!

After a short night with an incomplete puzzle the crew woke up with new energy to start their day. Still the puzzle fever caught most of the crew during breakfast…

Our chef-de-cuisine Hannah and so-chef Dave prepared biscuits and gravy which was a new experience for most of the crew and all of us loved it!

For lunch we had Kartoffelgratin (translated potato casserole) prepared by Maria before we prepared for our EVA!

Shortly after Nathan, Natalia, Ghanim and Maria stepped out to make their way to the Candor Chasma to collect rock samples we realized that the rovers Opportunity and Spirit had not been charging for the past 3 days.

This meant the crew had to troubleshoot and wait until the rovers were at full charge before we could use them. A second attempt was performed at 1430 and with fully charged rovers we made our way down the road. At the end of Galileo road, we realized that the weather was changing, and the EVA crew decided to return immediately. After this short EVA the crew was welcomed by the support crew and an amazing looking lava cake.

The afternoon was packed with crew members filling out the surveys for the leadership project, the continuous care of the ISRU plants (happily still alive) and continued work on the puzzle, which is still bugging us…. But we promise that we will not rest today until it is done!

Ad Astra!

Greenhab Report – February 20th

Crew 205 Green Hab Report – 20-FEB-2019

GreenHab Officer: Nathan Hadland

Environmental control: Heating and Cooling w/ ambient air (2.5 hrs)

Shade Cloth (40%)

Average temperatures:
Low: 17.3°C

High: 27.7°C

Hours of supplemental light: 5 Hours

Daily water usage for crops: 9 gallons

Water in Blue Tank: ~ 189 gallons

Crops watered at 0900 hours and 1800 hours:


1) 09.00 AM 19.8° C / 23 %
2) 12.00 PM 27.1° C / 20 %
3) 15:20 PM 19.3° C / 23 %
4) 18:00 PM 22.0° C / 30 %

Changes to crops: None.

Narrative: The nutrient supplementation in regolith experiments are still ongoing. However, due to mishaps in shipping, chlorophyll analysis may not be possible since the methanol has not arrived to the Hab (an organic solvent to help extract chlorophyll from the biomass). Consequently, these experiments may have to end without further analysis and they will be repeated upon return to Florida Institute of Technology using the same regolith samples. These experiments performed at the Hab do nevertheless represent the baseline tests astronauts will have to perform to determine the viability of substrates to support plant growth.

The nutrient deficiency observed in the tomato plants appear to be improving, since the yellow coloration is decreasing. The “Mystery Greens” from the spilled seed bank have sprouted and appear to be healthy, with two apparent variants of seedlings. Tomorrow, we will perform some additional grooming of dead biomass from several of the unhealthy plants to stimulate growth as well as contribute to our ISRU cellulosic ethanol study. Similar to yesterday, the door remained closed for most of the day due to the low ambient temperatures and overcast weather. However, at 1200 hours, the door was opened to allow for ambient airflow and closed at 1340 hours due to the drop in temperature.

Harvest: None.

Support/supplies needed: None.

Sol Summary – February 20th

Sol: 10

Summary Title: More Puzzling Weather on Mars

Author’s name: Hannah Blackburn

Mission Status: Nominal

Sol Activity Summary: We woke up to an overcast sky and biscuits with sausage gravy. The crew worked on a jigsaw puzzle of Mars while drinking their morning coffee. The EVA 14 team left the Hab at 1300 and quickly realized that Spirit and Opportunity did not fully charge. Atila reset the RAM’s external breaker which solved the problem. Approximately one hour later, the EVA 14 team set out again towards Candor Chasma on Spirit and Curiosity. However, when they had nearly reached their destination, they noticed fast-moving clouds and decided to return to the Hab to prevent weather-related issues. Later in the afternoon, light precipitation occurred while the crew ate chocolate lava cake and continued to piece together the puzzle. We finished the evening by enjoying movies from the Hab collection. Apollo 13 is a crew favorite.

Look Ahead Plan: If weather and ground conditions allow, although it currently seems unlikely, part of the crew will go on EVA 15 to The Moon. We will continue to work on developing Standard Operating Procedures for future crews.

Anomalies in work: Spirit and Opportunity did not fully charge. Resetting the RAM’s external breaker fixed the issue.

Weather: Overcast, light flurries starting at 1530

Crew Physical Status: Nominal


Reports to be filed: Sol Summary, Journalist Report, GreenHab Report, Operations Report, EVA report, EVA request

Support Requested: We did not receive any cloth towels with the initial supplies and therefore are on our last roll of paper towels. We would appreciate more paper or cloth towels for cleaning.

Copyright © The Mars Society. All rights reserved. | Main Site